OpenStudy (anonymous):

I need help solving T= 3U/e for U

OpenStudy (luigi0210):

Welcome to Openstudy!

OpenStudy (luigi0210):

First what you need to do is multiply both sides by e to get rid of the fraction.

OpenStudy (anonymous):

So I end up with e*T=3U

OpenStudy (luigi0210):

Correct, now do you know how to get rid of that 3?

OpenStudy (anonymous):

Nope, no clue

OpenStudy (luigi0210):

Well it's multiplication.. how do we undo multiplication?

OpenStudy (anonymous):

Division. Divide 3 by both sides?

OpenStudy (luigi0210):

Bingo

OpenStudy (anonymous):

Okay, it's 3/e*T=3U/3 and that would be 3/e*T=U?

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

Is that it?

OpenStudy (luigi0210):

Pretty much, simple right?

OpenStudy (anonymous):

Yeah, it was actually really simple. Thanks!

OpenStudy (luigi0210):

You are welcome :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."